Can you please help me. I'm want to solve this system of equation using the substitution method.
-3x+3y=4
-x+y=3

Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39626)

No intersection; parallel lines.

Answer by ikleyn(52862)

.
Can you please help me. I'm want to solve this system of equation using the substitution method.
-3x+3y=4
-x+y=3
~~~~~~~~~~~~~~~

-3x + 3y = 4,   (1)
-x  +  y = 3    (2)

From (2), express y = 3+x and substitute this expression into equation (1) replacing y. You will get

-3x + 3(3+x) = 4.

Simplify:

-3x + 9 + 3x = 4.

9 = 4.


Absurd ?  Sure.  It means that the original system has no solutions.
It is inconsistent. 


One can see it immediately/instantly after seeing the original system.
Indeed, multiply the second equation by 3. 
You will get an equivalent equation

-3x + 3y = 9.   (2')

Now compare (2') with (1). They have identical left sides, but different right sides.
Such a system is inconsistent.

For better understanding, see the lessons
    - Solution of the linear system of two equations in two unknowns by the Substitution method
    - Solution of the linear system of two equations in two unknowns using determinant
    - Solution of the linear system of two equations in two unknowns by the Elimination method
    - Geometric interpretation of the linear system of two equations in two unknowns
in this site.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Systems of two linear equations in two unknowns".